# Decoding Financial Health in Kenya’s Medical Insurance Sector: A Data-Driven Cluster Analysis

### v1

@@ 3.5   Time Series Clustering
Clustering time-series data is crucial for identifying trends and uncovering patterns across multiple entities [23]. In this study, we group insurance companies based on the latent representations $\mathbf{Z}\in\mathbb{R}^{N\times J\times 1}$ obtained from the LSTM model. These representations capture the temporal dynamics of financial ratios for each company while reducing the dimensionality of the data. Given the temporal nature of the data and potential structural variations, we employ the Dynamic Time Warping (DTW) algorithm [18], a robust method for measuring time-series similarity.
DTW is particularly suitable for datasets with temporal misalignments or missing values, as it aligns time-series data of unequal lengths [26]. Its insensitivity to outliers ensures that unusual data points do not disproportionately affect similarity measurements [39], making DTW ideal for clustering time-series data characterized by temporal shifts or distortions.
@@
Figure 4: DTW Distance Matrix Heatmaps. On the left is the unordered DTW distance matrix, which visualizes the pairwise distances between time series for all insurance companies without any specific order. On the right is the ordered DTW distance matrix, where the rows and columns are rearranged based on hierarchical clustering (complete linkage method). This reordering groups companies with similar time series patterns, allowing for a clearer identification of relationships and clusters within the data. The color intensity represents the magnitude of the DTW distance, with darker colors indicating smaller distances.
@@ 3.6   Evaluation of Clusters
The quality of the clusters identified in this analysis is evaluated using two widely recognized metrics: the Silhouette Score and the Elbow Method. These approaches provide complementary insights into the cohesiveness and separability of clusters and help determine the optimal number of clusters.
The Silhouette Score measures how well-separated the clusters are by assessing the similarity of each latent representation $z_{ij}$ to its assigned cluster compared to other clusters [9]. For a given latent value $z_{ij}$, the Silhouette Score $s_{ij}$ is calculated as:
@@
Figure 5: Cluster Evaluation Methods — Elbow and Silhouette Scores. The left plot shows the Silhouette Curve, indicating clustering quality, where higher values suggest better separation and cohesion of clusters. The right plot illustrates the Elbow Method, which identifies the optimal cluster number by the ”elbow” point where the rate of decrease in inertia slows.
@@ 4   Clustering Results
The LSTM architecture effectively compressed the high-dimensional dataset from $\mathbb{R}^{N\times J\times F}$ to a more concise representation in $\mathbb{R}^{N\times J\times F^{\prime}}$, preserving key temporal patterns essential for clustering and interpretation. By reducing the original 7-dimensional financial data vectors $\mathbf{r}_{ij}\in\mathbb{R}^{7}$ to 1-dimensional latent representations $z_{ij}\in\mathbb{R}^{1}$, the LSTM structure successfully captured meaningful temporal dependencies in the financial ratios of companies over time.
To improve interpretability, the financial ratios detailed in Table 2 were normalized prior to dimensionality reduction. Pairwise Dynamic Time Warping (DTW) distances were then computed between the latent representations $\mathbf{Z}$ of insurance companies and visualized as a distance matrix (Fig. 4, left). This matrix was reorganized using a dendrogram to enhance clarity, revealing the underlying cluster structures (Fig. 4, right). In the matrix, smaller blocks indicate tightly clustered companies with high similarity, while larger blocks reflect groups with greater dissimilarity. For instance, the dark-blue regions corresponding to MUA and Old Mutual form a distinct cluster, highlighting their strong similarity as shown in the reorganized matrix.
@@
The clustering results for $m=4$ are presented in Fig. 6, where the vertical axis represents the scaled latent time series data $\mathbf{Z}$ across 41 quarters for each company. Pairwise DTW distances between the latent time series were calculated, and the TimeSeriesKMeans algorithm grouped companies accordingly. Cluster centers, shown as bold lines, represent the average latent time series within each cluster, clearly highlighting overall trends. The latent time series of each company was overlaid with its respective cluster center to facilitate comparison. Each cluster is displayed in a separate subplot to improve clarity, enabling detailed observation of both individual company performance and broader group trends.
@@
Figure 6: Visual representation of time series clusters, with each cluster’s members and trends displayed in separate scatter plots. The top row contains the scatter plots for clusters 1 (top left) and 2 (top right). The bottom row displays clusters 3 (bottom left) and 4 (bottom right). The individual company-scaled time series (first feature) for each cluster are plotted with distinct colors, along with the smoothed cluster center line. The gradient fill highlights the cluster’s overall trend, and company names are annotated in the legend for easy identification. The x-axis represents time steps, while the y-axis corresponds to the values of the first reduced features for each company.
The four clusters identified in this study are defined as follows:
  • Cluster 1: Resilient Contributors – Companies demonstrating stable overall financial performance with occasional fluctuations.
  • Cluster 2: Emerging Contributors – Companies that began reporting data only in the later stages of the time series.
  • Cluster 3: Volatile Contributors – Companies exhibiting fluctuating financial performance with alternating high and low outcomes, along with occasional gaps in reporting.
  • Cluster 4: Early Contributors – Companies providing financial data primarily during the early phases of the time series, with sparse or missing data in later periods.
These clusters encapsulate diverse financial trajectories and reporting behaviors, offering a comprehensive perspective on the dynamics within the insurance sector. The distribution of companies within these clusters is detailed in Table 3.
CLUSTER INSURANCE COMPANIES SIZE CLASS NAME
1 AAR Insurance Kenya, APA Insurance, CIC General, First Assurance, GA Insurance, ICEA LION General, Kenindia, Madison, Pacis, Xplico 10 Resilient Contributors
2 MUA, Old Mutual, The Kenya Alliance 3 Emerging Contributors
@@
3 Allianz Insurance, Cannon General Insurance, Corporate Insurance, Sanlam, Takaful, Tausi, Trident 7 Volatile Contributors
@@
4 Britam General, Gateway, Heritage Insurance, Jubilee Health Insurance, Real Insurance, Resolution Health, Saham, UAP 8 Early Contributors
Table 3: Clustering results for insurance companies based on time series financial ratios data, utilizing KMeans and Dynamic Time Warping (DTW) methods. The table illustrates the distribution and classifying characteristics of each cluster.
@@ 5   Discussion
The clustering results derived from financial ratios and time series data provide significant insights into the operational and financial behaviors of insurance companies. Identifying four distinct clusters enables a nuanced understanding of variations in performance, reporting consistency, and systemic challenges within the industry.
Cluster 1: Resilient Contributors represents companies with predominantly strong financial stability, interrupted by sporadic declines in profitability. These fluctuations often reflect external shocks or market volatility, emphasizing the importance of robust risk management frameworks. This group, which includes most companies, aligns with broader industry trends that balance resilience with periodic vulnerabilities (Fig. 6, top right).
Cluster 2: Emerging Contributors comprise companies that began reporting data only in the latter part of the time series. Despite the early gaps, these companies have demonstrated strong financial performance in recent years, outperforming others in some metrics. This dynamic suggests strategic reentry into the market or adaptive responses to changing conditions, emphasizing improved data governance to facilitate accurate long-term evaluations (Fig. 6, bottom right).
Cluster 3: Volatile Contributors encompasses companies with inconsistent financial performance, characterized by fluctuating high and low values and occasional missing Gross Premium Income (GPI) data. These companies likely face systemic challenges, such as undercapitalization or operational inefficiencies, that require targeted regulatory scrutiny. Interestingly, some firms within this cluster exhibit relatively better data representation, suggesting that cluster outliers warrant tailored analysis to identify unique challenges (Fig. 6, top left).
Cluster 4: Early Contributors highlight issues related to data reporting irregularities. Companies in this group show consistent financial data only in the initial period, with missing data after 2020. This raises questions about their long-term operational continuity and viability, as abrupt data discontinuities may signify underlying financial or operational instability (Fig. 6, bottom left).
The clustering framework reveals that while Clusters 1 and 3 include companies with relatively stable data, Clusters 2 and 4 underscore the challenges of inconsistent reporting. For instance, companies in Cluster 3 showed improved data consistency up to 2020, after which reporting ceased. In contrast, Cluster 2 companies, despite their earlier absence, display high performance in later years, particularly in metrics like claims coverage, suggesting their suitability for specific insurance segments like medical coverage in recent periods.
These observations align with previous studies, such as [9] and [15], which highlight the utility of clustering methods in identifying homogeneous groups and the role of algorithm selection in shaping outcomes. The application of Dynamic Time Warping (DTW) in this study advances the analysis by capturing temporal patterns often overlooked in static assessments. However, the finding that Cluster 1’s stable companies also experience periodic losses challenges traditional assumptions of financial robustness, warranting further investigation into external drivers of these fluctuations.
Despite this analysis’s strengths, several limitations must be acknowledged. The capping of financial ratios, implemented to mitigate the influence of extreme values and outliers, may have led to the loss of valuable insights into companies with exceptionally high or low performance. Ratios such as the claims paid, claims incurred, expense, and underwriting profit ratios are particularly sensitive to this approach, potentially obscuring nuanced financial conditions. The analysis also excludes qualitative and non-financial factors, such as customer satisfaction or regulatory compliance, which could provide critical context. In addition, due to data constraints, reliance on common financial ratios excluded other essential measures, such as Return on Sales (ROS), Return on Assets (ROA), and Return on Equity (ROE). The study’s focus on the medical insurance sector further limits the generalizability of its findings to the broader insurance industry.
From a methodological perspective, the use of Long Short-Term Memory (LSTM) networks for dimensionality reduction and temporal pattern extraction proved effective in capturing meaningful insights from time-series financial data. The LSTM-based approach leverages its inherent ability to model sequential dependencies, making it particularly suitable for identifying clusters in datasets characterized by temporal variations and irregularities. Compared to traditional methods like Principal Component Analysis (PCA) or Gaussian Mixture Models (GMM), LSTMs offer the advantage of preserving temporal relationships, which are crucial in time-series analyses. Future studies could explore hybrid approaches, such as combining LSTMs with attention mechanisms, to enhance the interpretability of patterns extracted from the data.
@@
Data gaps present another challenge, particularly in clusters characterized by inconsistent reporting, such as ”Early Contributors” and ”Emerging Contributors.” Variations in reporting frequencies or missing data can distort temporal analyses and complicate cross-cluster comparisons. Addressing these inconsistencies through imputation techniques or enhanced data collection practices could improve the reliability and robustness of clustering outcomes.
To build on the findings of this study, future research should expand the analytical scope to include non-financial metrics, such as customer satisfaction, regulatory compliance, and competitive positioning. Incorporating granular data, such as company-level strategic decisions or macroeconomic indicators, could provide deeper insights into performance drivers. Additionally, exploring advanced machine learning models like Transformer networks or probabilistic clustering methods could reveal complementary perspectives and improve predictive accuracy. Addressing data inconsistencies in clusters with irregular reporting patterns, particularly Clusters 3 and 4, could result in actionable recommendations for enhancing governance and transparency in the sector.
By advancing the methodologies and addressing data challenges highlighted in this study, future research can contribute to a more comprehensive understanding of financial performance dynamics within the insurance industry. These efforts will support the development of robust, data-driven strategies for improving resilience and market transparency.

### v2

@@ 3.5   Time Series Clustering
Clustering time-series data is crucial for identifying trends and uncovering patterns across multiple entities [24]. In this study, we group insurance companies based on the latent representations $\mathbf{Z}\in\mathbb{R}^{N\times J\times 1}$ obtained from the LSTM model. These representations capture the temporal dynamics of financial ratios for each company while reducing the dimensionality of the data. Given the temporal nature of the data and potential structural variations, we employ the Dynamic Time Warping (DTW) algorithm [19], a robust method for measuring time-series similarity. DTW is particularly suitable for datasets with temporal misalignments or missing values, as it aligns time-series data of unequal lengths [27]. Its insensitivity to outliers ensures that unusual data points do not disproportionately affect similarity measurements [41], making DTW ideal for clustering time-series data characterized by temporal shifts or distortions.
@@
Figure 4: On the left is the unordered DTW distance matrix, which visualizes the pairwise distances between time series for all insurance companies without any specific order. On the right is the ordered DTW distance matrix, where the rows and columns are rearranged based on hierarchical clustering (complete linkage method). This reordering groups companies with similar time series patterns, allowing for a clearer identification of relationships and clusters within the data. The color intensity represents the magnitude of the DTW distance, with darker colors indicating smaller distances.
@@ 3.6   Evaluation of Clusters
The quality of the clusters identified in this analysis is evaluated using two widely recognized metrics: the Silhouette Score and the Elbow Method. These approaches provide complementary insights into the cohesiveness and separability of clusters and help determine the optimal number of clusters. The Silhouette Score measures how well-separated the clusters are by assessing the similarity of each latent representation $z_{ij}$ to its assigned cluster compared to other clusters [10]. For a given latent value $z_{ij}$, the Silhouette Score $s_{ij}$ is calculated as:
@@
Figure 5: The left plot shows the Silhouette Curve, indicating clustering quality, where higher values suggest better separation and cohesion of clusters. The right plot illustrates the Elbow Method, which identifies the optimal cluster number by the ”elbow” point where the rate of decrease in inertia slows.
3.7   Time Series Financial Performance Analysis
The time-series analysis of financial performance provides crucial insights into the trends and dynamics shaping the Kenyan medical insurance industry. The five heatmaps illustrate variations in key financial indicators, including market share, net earned premium income, underwriting profit, the number of new policies, and total policies issued over the years. The heatmaps present the trends across different insurance companies from 2013 to 2023.
The heatmap 6 illustrates the distribution of market share among various insurers over time. The visualization provides insights into the competitive landscape of the insurance industry, highlighting both dominant players and those with fluctuating or marginal market shares. One of the most notable observations is the consistent dominance of Jubilee Health, which maintains the highest market share across multiple time periods. Other significant players, such as AAR and APA Insurance also hold substantial market shares in different periods. These companies appear to have maintained their competitive positioning over time. In contrast, several insurers experience noticeable fluctuations in their market share. Resolution Health, UAP, and Old Mutual demonstrate periods of growth, where they hold a larger portion of the market, followed by stabilization or decline. This suggests that while some firms experience temporary surges, maintaining a dominant position in the insurance sector requires sustained strategic efforts. Additionally, some insurers, such as CIC General and Britam General, maintain relatively stable but moderate market shares, indicating steady performance without dramatic shifts.
Figure 6: Heatmaps illustrating the time series of the market share distribution among insurers. The color intensity represents the proportion of the total market share at each time period, with higher values indicated by red shades. At each time point, the market shares sum to 100%.
Net earned premium income, shown in Figure 7 (top left), represents the portion of premiums retained after deductions for reinsurance and adjustments for policy cancellations. A closer examination of the heatmap reveals that while some companies consistently report stable or increasing premium income, others experience significant fluctuations. Companies such as Takaful and Tausi demonstrate low but steady growth in premium income, indicating robust underwriting practices and customer retention. Conversely, firms such as Madison Insurance exhibit sporadic premium inflows, suggesting challenges in policy sales, customer acquisition, or retention.
The underwriting profit heatmap, depicted in Figure 7 (top right), highlights profitability trends across insurance firms. Companies like Jubilee Health and Heritage Insurance consistently show strong underwriting profits, reflecting effective risk management and operational efficiency. However, some insurers, including CIC General and Resolution Health, report periods of negative underwriting profit, suggesting high claims payouts or excessive operational costs. Periodic losses for multiple companies emphasize industry-wide challenges in maintaining sustainable underwriting margins, likely due to fluctuating claims and expense ratios.
The number of new policies heatmap, illustrated in Figure 7 (bottom left), provides insights into policy issuance trends across different insurers. A significant observation is that Britam General records a substantial increase in new policies in recent years, positioning it as a leading player in customer acquisition. However, firms such as MUA, Takaful, and Pacis report minimal new policy sales, suggesting difficulties in attracting new customers. These disparities in new policy issuance highlight differences in market penetration strategies and product offerings among insurers.
The total policies heatmap, as shown in Figure 7 (bottom right), provides insights into the overall size of an insurer’s customer base. Britam General maintains a substantial policyholder base, reflecting strong customer retention and renewal strategies. Conversely, insurers like AAR Insurance, Cannon Insurance, and MUA show relatively low total policy counts, suggesting weaker customer retention or lower demand for their insurance products. The presence of abrupt changes in policy numbers for some firms, such as Xplico, could indicate strategic shifts, acquisitions, or product restructuring.
Figure 7: Heatmaps illustrating key financial and operational metrics of various insurance companies over time. (Top Left) Net Earned Premium Income trends from 2013 to 2023. (Top Right) Underwriting Profit patterns over the same period. (Bottom Left) Number of New Policies issued each year for the general insurance. (Bottom Right) Total Policies in force over time. The color intensity represents the magnitude of each metric, with higher values indicated by red shades.
@@ 4   Clustering Results
The LSTM architecture effectively compressed the high-dimensional dataset from $\mathbb{R}^{N\times J\times F}$ to a more concise representation in $\mathbb{R}^{N\times J\times F^{\prime}}$, preserving key temporal patterns essential for clustering and interpretation. By reducing the original 7-dimensional financial data vectors $\mathbf{r}_{ij}\in\mathbb{R}^{7}$ to 1-dimensional latent representations $z_{ij}\in\mathbb{R}^{1}$, the LSTM structure successfully captured meaningful temporal dependencies in the financial ratios of companies over time. To improve interpretability, the financial ratios detailed in Table 2 were normalized prior to dimensionality reduction. Pairwise Dynamic Time Warping (DTW) distances were then computed between the latent representations $\mathbf{Z}$ of insurance companies and visualized as a distance matrix (Fig. 4, left). This matrix was reorganized using a dendrogram to enhance clarity, revealing the underlying cluster structures (Fig. 4, right). In the matrix, smaller blocks indicate tightly clustered companies with high similarity, while larger blocks reflect groups with greater dissimilarity. For instance, the darker regions corresponding to MUA and Old Mutual form a distinct cluster, highlighting their strong similarity as shown in the reorganized matrix.
@@
4.1   K-Means Clustering
K-Means clustering is a widely used unsupervised machine learning technique that partitions data into $m$ distinct clusters based on similarity measures. In this study, the TimeSeriesK-Means algorithm was applied to group insurance companies according to the temporal patterns in their financial performance. Unlike traditional K-Means, which relies on Euclidean distance, TimeSeriesK-Means incorporates Dynamic Time Warping (DTW) to account for temporal misalignments in financial data. In this study, we calculated Pairwise DTW distances between the latent time series, and the TimeSeriesK-Means algorithm grouped companies accordingly. The clustering results for $m=4$ are presented in Fig. 8, where the vertical axis represents the scaled latent time series data $\mathbf{Z}$ across 41 quarters for each company.
@@ 4.1   K-Means Clustering
Figure 8: The figure presents scatter plots of the scaled latent time series for the four identified clusters. The top-left and top-right plots correspond to Clusters 1 and 2, respectively, while the bottom-left and bottom-right plots depict Clusters 3 and 4. Each company’s time series is color-coded within its respective cluster, with the bold lines representing cluster centers, which indicate the average latent time series trend. The gradient fill emphasizes overall cluster trends, while company names are annotated in the legend for easy identification. The x-axis represents time steps, and the y-axis denotes the first reduced feature values for each company. Cluster compositions include seven companies in Cluster 1, ten in Cluster 2, eight in Cluster 3, and three in Cluster 4.
Cluster 1 comprises insurance companies that exhibit inconsistent financial performance, as reflected in the time series trends across various financial indicators. These companies show frequent fluctuations in their values, with periods of both high and low performance, as well as occasional gaps in their Gross Premium Income data. The cluster highlights erratic movements for several insurers, characterized by extreme peaks and troughs, indicating instability in their premium collections and overall financial performance. The presence of missing data points further underscores the possibility of financial distress or reporting inconsistencies, which may require regulatory intervention or targeted support. These companies generally hold low and fluctuating market shares, reinforcing their lack of stable positioning within the industry (see Fig. 7, top left). Moreover, in the Net Earned Premium Income heatmap, some of these insurers display inconsistent earnings, with sporadic bright patches indicating occasional financial improvements. However, these improvements do not appear to be sustained over time, aligning with underwriting profit trends that also suggest volatility. The underwriting profit heatmap reveals that some of these firms experience periods of negative profits, signaling poor risk management or high claims ratios that impact financial stability. Interestingly, the new policies and total policies heatmaps provide additional insights into the performance of these firms. Many of these companies exhibit low or inconsistent new policy issuance, indicating weak customer acquisition strategies or an inability to expand their policyholder base. Some firms may experience occasional spikes in new policies, but these do not translate into sustained growth in total policies over time (see Fig. 7). As shown in Fig. 6, these companies also struggle to maintain a consistent market share, which may indicate challenges related to customer acquisition, retention, or competitive pricing strategies.
Cluster 2 consists of insurers with relatively stable financial performance, punctuated by occasional fluctuations in profitability (Fig. 3.5, top right). These companies generally exhibit strong financial resilience, managing to maintain consistent market positions despite periodic declines. The cluster includes a group of insurers such as AAR Insurance, APA Insurance, CIC General, ICEA LION, First Assurance, GA Insurance, Kenindia, Madison, Pacis, and Takaful, which navigate financial cycles with intermittent periods of decline. However, compared to firms in Cluster 1, these insurers demonstrate stronger recoveries and a general tendency to stabilize after downturns. The fluctuations observed in this cluster are less extreme than those in Cluster 1, indicating a more structured approach to financial and operational management. The market share heatmap (Fig. 6) reveals that many Cluster 2 insurers hold moderate to significant portions of the market share, reflecting their relatively stable competitive positioning. Companies such as ICEA LION and APA Insurance maintain consistent market shares, while others, like CIC General and First Assurance, show steady but moderate market penetration. Their ability to sustain their positions over time suggests strong customer trust, effective underwriting strategies, and solid financial management. Further insights can be drawn from the Net Earned Premium Income and Underwriting Profit heatmaps, where these firms demonstrate better financial stability than Cluster 1 companies. While sporadic downturns are observed, they do not persist for extended periods, suggesting that these firms have diversified revenue streams, efficient claims management, or effective cost-control measures that mitigate prolonged losses. The new policies and total policies heatmaps further support this classification, showing consistent issuance of new policies for most companies in this cluster. The ability to maintain a steady flow of new policies indicates market confidence and customer retention strategies that contribute to their financial resilience (see Fig. 7).
Cluster 3 comprises companies that were once dominant players in the insurance industry but have experienced a transition in their financial performance over time (see Fig. 3.5, bottom left). Contrary to earlier assumptions of data discontinuity, the time series trends, market share, and financial indicators reveal that these companies remain active, though their market positioning has evolved. Notably, firms such as Jubilee Health, Britam General, UAP, and Resolution Health have maintained substantial market presence and financial strength, even though their trajectories have fluctuated over time. The market share heatmap (Fig. 6) confirms this trend, showing that Jubilee Health, Britam General, and UAP have consistently held significant portions of the market over time. While their market shares may have fluctuated, they remain among the industry leaders, demonstrating resilience amid competitive pressures and economic shifts. Some companies in this cluster, such as Resolution Health, initially exhibited strong performance but later experienced gradual market share reductions, indicating possible strategic shifts, increased competition, or structural changes within the industry. Furthermore, the total policies and new policies heatmaps indicate that these insurers have maintained a steady flow of policy acquisitions over time. The Net Earned Premium Income heatmap shows that Jubilee Health, Britam General, and UAP continue to generate high premium revenues, reflecting strong underwriting performance and customer retention. The underwriting profit heatmap highlights some variability, with periods of high profitability followed by occasional dips. This suggests that while these companies have managed risk effectively, external shocks, regulatory changes, or claims volatility have impacted profitability at certain intervals (see Fig. 7).
Cluster 4 (Fig. 3.5, bottom right) represents a group of insurers that only began reporting financial data in later years but have since demonstrated strong financial performance and market expansion. These companies, including MUA, Old Mutual, and The Kenya Alliance, initially showed minimal or no data activity but have experienced rapid growth, outperforming other clusters in certain financial indicators. The market share heatmap (Fig. 6) supports this observation, as MUA, Old Mutual, and The Kenya Alliance show significant increases in market share in recent years. Unlike insurers in Cluster 3, which have gradually lost market share over time, these firms have expanded aggressively, capturing a larger portion of the industry. Additionally, the total policies and new policies heatmaps highlight a surge in policy issuance, confirming that these companies are actively acquiring new customers. Financially, Cluster 4 companies exhibit progressive improvements in premium income and underwriting profits. The Net Earned Premium Income heatmap shows a steady increase, indicating that these firms have successfully attracted policyholders and generated significant revenue from premiums. The underwriting profit heatmap reveals that their profitability has grown alongside premium earnings, suggesting effective risk management and claims control mechanisms (see Fig. 7).
4.2   Hierarchical Clustering
Hierarchical clustering was performed to analyze the structural relationships between insurance companies based on their financial time-series data. Unlike K-Means, which partitions the data into a predefined number of clusters, hierarchical clustering organizes companies into a hierarchical structure, allowing for a more flexible and interpretable grouping process [24]. The hierarchical clustering method employed a precomputed Dynamic Time Warping (DTW) distance matrix to capture temporal dependencies in financial ratios. Using the complete linkage method, which minimizes the maximum inter-cluster distance during merging, the resulting dendrogram (Fig. 9) reveals well-separated groups with companies exhibiting high intra-cluster similarity.
A key advantage of hierarchical clustering is its ability to validate the strong relationships between closely connected companies. Unlike K-Means, which can be sensitive to initialization and predefined cluster counts, hierarchical clustering naturally groups companies with highly similar financial trajectories, as evidenced by the tightly clustered firms in the dendrogram. The dendrogram’s vertical axis represents cluster distances, illustrating which companies are most similar and which groups maintain a degree of separation. Firms that merge at lower distance thresholds exhibit near-identical financial behavior, making hierarchical clustering a valuable tool for confirming firm-level similarities with higher granularity than K-Means. For example, MUA and Old Mutual form a distinct, compact subgroup, reinforcing their strong financial resemblance. Similarly, Cannon Insurance and Corporate Insurance in Cluster 3 remain closely linked, confirming their shared market dynamics and financial stability over time. The four-cluster partitioning aligns well with K-Means results while offering a more nuanced perspective on firm relationships. The hierarchical approach validates the grouping of insurance companies, reinforcing the robustness of the clustering methodology while allowing for a deeper exploration of industry dynamics.
Figure 9: Hierarchical clustering dendrogram of insurance companies based on their financial time series representations. The vertical axis represents cluster distance, illustrating the hierarchical relationships among companies. The four main clusters identified in the analysis are distinguished by different colors: Cluster 1 with three companies, Cluster 2 with three companies, Cluster 3 with eighteen companies, and Cluster 4 with two companies.
The distribution of companies within these clusters identified by hierarchical clustering and K-Means is detailed in Table 3.
CLUSTER K-Means Clusters Hierarchical Clusters
@@ 4.2   Hierarchical Clustering
1 Allianz Insurance, Cannon Insurance, Corporate Insurance, Sanlam, Takaful, Tausi, Trident Takaful, Sanlam, Trident
2 AAR Insurance Kenya, APA Insurance, CIC General, First Assurance, GA Insurance, ICEA LION, Kenindia, Madison, Pacis, Xplico Xplico, Allianz Insurance, Tausi, The Kenya Alliance
@@
3 Britam General, Gateway, Heritage Insurance, Jubilee Health, Real Insurance, Resolution Health, Saham, UAP UAP, Jubilee Health, Resolution Health, Saham, Cannon Insurance, Corporate Insurance, AAR Insurance, Britam General, First Assurance, Pacis, ICEA LION, Madison, APA Insurance, GA Insurance, CIC General, Kenindia, Gateway, Real Insurance, Heritage Insurance
4 MUA, Old Mutual, The Kenya Alliance MUA, Old Mutual
Table 3: Comparison of insurance company assignments between K-Means and Hierarchical Clustering based on financial time series data. The table highlights similarities and differences in how each method grouped companies, emphasizing the closer relationships captured by hierarchical clustering.
@@ 5   Discussion
The clustering results derived from financial ratios and time series data provide significant insights into the operational and financial behaviors of insurance companies. Identifying four distinct clusters enables a nuanced understanding of variations in performance, reporting consistency, and systemic challenges within the industry. The results indicate that certain insurers maintain financial stability and steady growth, while others exhibit volatility or inconsistent reporting practices, posing challenges to risk assessment and market transparency.
The implications of these findings extend to regulatory authorities, particularly the Kenya Insurance Regulatory Authority (IRA). Given the inconsistencies observed in Clusters 3 and 4, regulatory interventions could be implemented to enforce more stringent reporting standards. Mandatory reporting frameworks would ensure that all insurance companies submit timely and accurate financial data, reducing discrepancies and improving market transparency. Standardizing financial disclosures, particularly in underwriting profit and premium income, would allow for better comparability across insurers, facilitating more informed decision-making for both regulators and policyholders. Inconsistent reporting not only undermines risk assessment but also distorts industry performance metrics, which could lead to misinformed policy interventions. Enhanced regulatory oversight would mitigate these issues, ensuring that the insurance sector remains robust and accountable.
The clustering results also provide valuable insights for investors and policyholders in making informed decisions. A key takeaway from the analysis is that insurers can be categorized based on financial stability, growth trends, and risk exposure. Introducing a stability rating system could significantly enhance consumer confidence by classifying insurers based on underwriting profitability, policy retention rates, and premium consistency. Such a classification system would help policyholders identify insurers that are financially stable versus those exhibiting volatility, enabling better selection of insurance providers. Investors, on the other hand, could use these insights to assess risk-adjusted returns and make strategic investment decisions. This approach mirrors existing credit rating frameworks and could be adopted within the insurance sector to improve market efficiency. The classification of the companies aids stakeholders in identifying stable and high-risk insurers, making informed market decisions, and formulating risk mitigation strategies.
To ensure the robustness of the clustering methodology, hierarchical clustering was introduced alongside K-Means. While the clustering assignments remained largely consistent, hierarchical clustering provided further validation of the time-series-based groupings. Future work could incorporate additional cross-validation techniques, such as bootstrapping or reassignment consistency tests, to further confirm cluster stability. The application of Dynamic Time Warping (DTW) in this study is particularly suited to capturing temporal financial patterns, as it preserves the sequential dependencies in the data, distinguishing insurers with consistent long-term performance from those exhibiting sporadic fluctuations.
To enhance the contextual understanding of the clustered companies, additional firm-level attributes were considered, including market share, total policyholder base, and financial strength indicators such as underwriting profit trends. These attributes help explain variations in clustering results, as larger firms with significant market share tend to exhibit stable financial trends, whereas smaller firms or those with irregular policy issuances are more prone to volatility. The inclusion of this data provides a more holistic view of each insurer’s market positioning. Macroeconomic variables, such as inflation, interest rates, and regulatory changes, can significantly impact an insurer’s financial performance. For instance, rising inflation can increase claims costs, thereby affecting underwriting profits. Similarly, changes in regulatory requirements, such as solvency mandates, can influence capital adequacy ratios, impacting insurer clustering. The results indicate that insurers in different clusters may respond differently to economic shocks; for example, companies in Cluster 2 may exhibit resilience, whereas those in Cluster 1 may face greater financial distress. Future research could integrate macroeconomic indicators to further analyze these relationships.
This study contributes to both theory and practice by leveraging DTW-based clustering, which offers a novel approach to analyzing financial time-series data within the insurance sector. Compared to conventional clustering methods such as PCA or Gaussian Mixture Models, DTW captures temporal dependencies more effectively, leading to more accurate segmentation of insurers. These insights align with previous studies that emphasize the role of time-series analysis in financial modeling. From a managerial perspective, the results provide actionable insights. Companies in Cluster 1 could benefit from improving operational efficiency and risk management to stabilize their financial performance. Regulators could prioritize firms in Clusters 3 and 4, enforcing stricter reporting requirements to enhance data consistency. Insurers within Cluster 2 can leverage their financial stability to expand market share and introduce innovative insurance products.
Despite this analysis’s strengths, several limitations must be acknowledged. The analysis excludes qualitative and non-financial factors, such as customer satisfaction or regulatory compliance, which could provide critical context. In addition, due to data constraints, reliance on common financial ratios excluded other essential measures, such as Return on Sales (ROS), Return on Assets (ROA), and Return on Equity (ROE). The study’s focus on the medical insurance sector further limits the generalizability of its findings to the broader insurance industry. From a methodological perspective, the use of LSTM networks for dimensionality reduction and temporal pattern extraction proved effective in capturing meaningful insights from time-series financial data. However, while LSTM excel at modeling sequential dependencies and preserving temporal relationships, their technique can limit interpretability compared to traditional methods such as Principal Component Analysis (PCA) or Gaussian Mixture Models (GMM). Additionally, LSTM require large datasets for optimal training, making them sensitive to data availability and quality. Future studies could explore hybrid approaches, such as integrating LSTM with attention mechanisms, to enhance interpretability and improve performance in cases with limited financial data.
@@
Data gaps present another challenge, particularly in clusters characterized by inconsistent reporting. Variations in reporting frequencies or missing data can distort temporal analyses and complicate cross-cluster comparisons. Addressing these inconsistencies through imputation techniques or enhanced data collection practices could improve the reliability and robustness of clustering outcomes. To build on the findings of this study, future research should expand the analytical scope to include non-financial metrics, such as customer satisfaction, regulatory compliance, and competitive positioning. Incorporating granular data, such as company-level strategic decisions or macroeconomic indicators, could provide deeper insights into performance drivers. Additionally, exploring advanced machine learning models like probabilistic clustering methods could reveal complementary perspectives and improve predictive accuracy. Addressing data inconsistencies in clusters with irregular reporting patterns, particularly Clusters 3 and 4, could result in actionable recommendations for enhancing governance and transparency in the sector. By advancing the methodologies and addressing data challenges highlighted in this study, future research can contribute to a more comprehensive understanding of financial performance dynamics within the insurance industry. These efforts will support the development of robust, data-driven strategies for improving resilience and market transparency.